Let 1 0 0 0 0 0 2 0 0 0 A=0 0 3 0 0 0 0 0 4 0 0 0 0 0 5 (a) (5 points) Find the general solution to the sysem of equations I(b) (5 points) Solve the initial value problem x1(0) = x2(0) - = ... = C5(0) = 5.
Answer #1

Krest, 02 oowoo OoOJO ooooh O ny System, n - 2 o o zoo ooyo n3 vy us -5 - 202 2 (3 n = 404 -6 ns =50 24 = 494 → -kyet Solutio